OpenStudy (anonymous):

A line passes through the points (-8, 7) and (x2, y2), and has a slope of -1/10. If point (x2, y2) is located in quadrant I, find x2.

OpenStudy (aravindg):

slope= \(\dfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

i know that i just dont know how to find x2

OpenStudy (aravindg):

ok can you show how much you have progressed?

OpenStudy (anonymous):

thats how i think i can do it but i dont if im right im clueless how to do this problem y2-7 (over) x2- -8= -1/10

OpenStudy (aravindg):

ok ?

OpenStudy (aravindg):

proceed

OpenStudy (anonymous):



OpenStudy (anonymous):

is x2 2

OpenStudy (anonymous):

and y1 is 6

OpenStudy (anonymous):

so am i right

OpenStudy (anonymous):